# Words Like Knives: Backstory-Personalized Modeling and Detection of Violent Communication

### v1

@@ Abstract
Conversational breakdowns in close relationships are deeply shaped by personal histories and emotional context, yet most NLP research treats conflict detection as a general task, overlooking the relational dynamics that influence how messages are perceived. In this work, we leverage nonviolent communication (NVC) theory to evaluate LLMs in detecting conversational breakdowns and assessing how relationship backstory influences both human and model perception of conflicts. Given the sensitivity and scarcity of real-world datasets featuring conflict between familiar social partners with rich personal backstories, we contribute the PersonaConflicts Corpus [^1] Dataset link will be made available upon publication, a dataset of $N=5,772$ naturalistic simulated dialogues spanning diverse conflict scenarios between friends, family members, and romantic partners. Through a controlled human study, we annotate a subset of dialogues and obtain fine-grained labels of communication breakdown types on individual turns, and assess the impact of backstory on human and model perception of conflict in conversation. We find that the polarity of relationship backstories significantly shifted human perception of communication breakdowns and impressions of the social partners, yet models struggle to meaningfully leverage those backstories in the detection task. Additionally, we find that models consistently overestimate how positively a message will make a listener feel. Our findings underscore the critical role of personalization to relationship contexts in enabling LLMs to serve as effective mediators in human communication for authentic connection.
@@ 1 Introduction
Human communication is inherently contextual, especially in close relationships such as those between romantic partners, family members, or friends. In these settings, speakers and listeners share a rich history and tailor their messages based on past experiences, personal sensitivities, and relational dynamics Isaacs and Clark (1987); Wheatley et al. (2023); Pillemer (1992). It is also in these intimate contexts where conversational breakdowns are most likely to occur—moments when language evokes hurt, misunderstanding, or conflict. Crucially, whether or not a message constitutes a breakdown depends on how it is interpreted in light of the speaker-listener relationship Zhou et al. (2023b); Schurz et al. (2021); Dvash and Shamay-Tsoory (2014). For example, as shown in Figure 1, a mother telling her daughter “You should spend more time at home with me” may be perceived as manipulative and controlling in one context (Backstory B), or as a tender expression of grief in another (Backstory A).
@@
To address these gaps, we develop a framework that simulates and analyzes communication breakdowns in intimate conversations, taking into account the relationship context. Our approach draws on Nonviolent Communication (NVC) theory Rosenberg and Chopra (2015), a structured approach widely used in conflict resolution and therapeutic settings, to guide our evaluation of LLM’s detecting harmful communication. First, we introduce PersonaConflicts Corpus, a dataset of $N=5{,}772$ simulated conflict dialogues between familiar social partners, spanning across diverse scenarios. For each conversation, we generate two distinct backstories: a positive backstory, which frames one character’s actions as more understandable or sympathetic, and a negative backstory, which portrays the same character in a more problematic or blameworthy light Moon et al. (2024). Through human validation, we show that scenarios and backstories are largely believable.
@@ 2.1 Conversational Breakdown Detection and Reframing
In the area of AI-mediated communication (AIMC), text rewriting can be used to improve interpersonal outcomes like empathy or social connection by suggesting changes to the tone or style of a message at the right time Hancock et al. (2020). To support such systems, prior tasks propose detecting conversational breakdowns between people, and reframing messages to be more empathetic.
@@ 4 PersonaConflicts Corpus
We introduce the PersonaConflicts Corpus, a dataset of realistic conflict and non-conflict scenarios simulated using LLMs (Figure 3). Given the sensitivity of conflict-laden conversations between familiar social partners, no prior open-source datasets are suitable to observe violent communication breakdowns in dialogue. As such, we draw on recent work in backstory generation and persona alignment Moon et al. (2024); Jiang et al. (2024); Hu and Collier (2024) as well as multi-agent simulation Zhou et al. (2023a, 2025); Ahmad et al. (2025); Kim et al. (2024) to generate conflict-laden scenarios which inject violent communication practices in turns as well as generating a set of non-conflict conversations. Our simulation setup uses the gpt-4 model and all prompts are included in Appendix A. We later discuss our human validation to evaluate soundness of our dataset (Section 5).
@@ 5 Human Study and Annotation
To answer RQ1, how backstory influences people’s perceptions of conflict, we conducted a human study to assess the impact of backstory variant on perception of conflict at the conversation level and turn level, as well as to evaluate the quality of our dataset and obtain fine-grained annotations of violent or non-violent communication types.
@@
We conducted a between-subjects study where participants are assigned to a positive or negative backstory. First, participants read the background of the characters and the conversation and rated overall measures of the conversation. Then, they were asked to rate each turn of the dialogue (See Figure 4 for an example of our user interface). The average work time was 17.28 minutes, and workers were paid \$3 for each HIT. Two independent workers completed each HIT for inter-annotator agreement calculation, resulting in a total of 480 annotations (3,474 turns rated across 120 conversations with 2 versions of backstory and 2 annotators per conversation). All annotation templates are included in Appendix B.
@@
NEG Turn is problematic (4 point) .$80$44.44.44.44
Emotional impact (3 point) .$79$46.46.46.46
POS Turn is problematic (4 point) .$78$34.34.34.34
Emotional impact (3 point) .$78$42.42.42.42
@@ 6 Effect of Backstory Personalization on Human Participants
As shown in Figure 6, we found that providing relationship backstory significantly shifted participant perceptions of communication quality. Specifically, for negative backstories, characters were rated as more problematic both at the turn level ($t(1083)=3.73$, $p=0.0002$, Cohen’s $d=0.23$) and at the overall conversation level ($t(232)=4.18$, $p<0.0001$, Cohen’s $d=0.55$). Additionally, with negative backstories, participants found the character’s communication less understandable ($t(232)=-4.70$, $p<0.0001$, Cohen’s $d=-0.61$) and interpreted their behavior as expressing more negative intent towards the other character ($t(232)=-4.95$, $p<0.0001$, Cohen’s $d=-0.65$). Notably, participants expressed significantly lower sympathy towards the character when a negative backstory was present ($t(232)=-7.05$, $p<0.0001$, Cohen’s $d=-0.92$), indicating a strong effect of relationship context on social judgments. These findings demonstrate that backstory personalization meaningfully influences how human raters interpret conflict.
@@
Table 2: F1 scores (in percentage) for predicting turn problematicness and emotional impact across models, conditions, and backstory types. Bold indicates the highest score in a model/backstory group; underline indicates the second highest. Significance stars denote statistical difference from the prior condition: * p$<$0.05, ** p$<$0.01, *** p$<$0.001.
@@ 7.2 Results and Discussion
These results suggest that while GPT-4o is the most consistent with human perception across both tasks, LLaMA is prone to strong overestimation in both problematic detection and emotional response. Gemini’s behavior is more sensitive to backstory polarity, with notable shifts in prediction direction depending on whether a speaker is portrayed sympathetically or not, however these shifts are more extreme than human annotators’ ratings. Overall, our findings indicate that models might not be effectively leveraging relationship backstories to tailor understanding of conversational dynamics, in alignment with human perception.
@@ 8 Conclusion
First, while our dataset and evaluation pipeline offer a scalable and theory-informed way to study communication breakdowns, our conversations are generated via simulation, which may not capture the full complexity of real-world interpersonal dynamics Wang et al. (2025). Language models can mirror patterns of speech and emotion, but they lack lived experience and embodied context. Furthermore, our work is limited to a single modality, whereas conflict can also be interpreted through multimodal channels, such as through pitch and tone Tschacher et al. (2014). Thus, findings from our simulated dialogue corpus may not fully generalize to naturally occurring conflicts between real people, where nonverbal cues, power dynamics, and history shape interpretation in more nuanced ways.
However, consistent with prior findings, our generated conversations were deemed largely naturalistic Bao et al. (2023); Chuang et al. (2024); Hu and Collier (2024) and obtaining datasets of conflict-laden conversation between loved ones is extremely challenging to obtain while protecting user privacy. Our work addresses a step towards the aim of relationship contextualization in conflict understanding via language models, and future works can improve upon bridging the gap between simulation and real-world settings through methodological improvements Finch and Choi (2024) as well as evaluating how relationship-informed mediation tools perform in live settings with human users.
Second, we did not analyze every possible outcome measure collected during the human study. Our focus was primarily on a subset of key social-interpretive variables such as problematicness, sympathy, and intention, but future work could explore additional dimensions such as emotional volatility, perceived agency, or trust.
@@
All studies conducted in this work were classified under Institutional Review Board (IRB) exemption status. While our work aims to enhance interpersonal understanding and mitigate conflict through AI-mediated communication, the use of simulated dialogues and backstories about emotionally sensitive relationships—such as those between romantic partners or family members—raises concerns around realism and potential misuse. Although we do not collect or model real personal data, generated dialogues might still resemble real-life situations and emotional dynamics. If deployed in real-world applications, such systems could be used to influence perceptions of others, shape interpretations of interpersonal interactions, or even manipulate emotional outcomes, especially in high-stakes or abusive relationships. It is crucial that such tools remain assistive rather than prescriptive, providing support while preserving user autonomy and avoiding overreach in delicate relational contexts.
@@ Acknowledgements
We would like to thank all of our participants and teammates for their invaluable contributions to this project. Special thanks to Ashish Sharma for feedback on the project. This work was funded in part by DSO National Laboratories and supported by the Defense Advanced Research Projects Agency (DARPA) under Agreement No. HR00112490410.
@@ A.5 Model Experiment Prompts
Appendix B Annotation Templates
Appendix C Supplementary Results

### v2

@@ Abstract
Conversational breakdowns in close relationships are deeply shaped by personal histories and emotional context, yet most NLP research treats conflict detection as a general task, overlooking the relational dynamics that influence how messages are perceived. In this work, we leverage nonviolent communication (NVC) theory to evaluate LLMs in detecting conversational breakdowns and assessing how relationship backstory influences both human and model perception of conflicts. Given the sensitivity and scarcity of real-world datasets featuring conflict between familiar social partners with rich personal backstories, we contribute the PersonaConflicts Corpus [^1] Dataset available at https://github.com/mitmedialab/persona-conflicts-corpus-emnlp-2025, a dataset of $N=5,772$ naturalistic simulated dialogues spanning diverse conflict scenarios between friends, family members, and romantic partners. Through a controlled human study, we annotate a subset of dialogues and obtain fine-grained labels of communication breakdown types on individual turns, and assess the impact of backstory on human and model perception of conflict in conversation. We find that the polarity of relationship backstories significantly shifted human perception of communication breakdowns and impressions of the social partners, yet models struggle to meaningfully leverage those backstories in the detection task. Additionally, we find that models consistently overestimate how positively a message will make a listener feel. Our findings underscore the critical role of personalization to relationship contexts in enabling LLMs to serve as effective mediators in human communication for authentic connection.
@@ 1 Introduction
Human communication is inherently contextual, especially in close relationships such as those between romantic partners, family members, or friends. In these settings, speakers and listeners share a rich history and tailor their messages based on past experiences, personal sensitivities, and relational dynamics Isaacs and Clark (1987); Wheatley et al. (2023); Pillemer (1992). It is also in these intimate contexts where conversational breakdowns are most likely to occur—moments when language evokes hurt, misunderstanding, or conflict. Crucially, whether or not a message constitutes a breakdown depends on how it is interpreted in light of the dyadic relationship Zhou et al. (2023b); Schurz et al. (2021); Dvash and Shamay-Tsoory (2014). For example, as shown in Figure 1, a mother telling her daughter “You should spend more time at home with me” may be perceived as manipulative and controlling in one context (Backstory B), or as a tender expression of grief in another (Backstory A).
@@
To address these gaps, we develop a framework that simulates and analyzes communication breakdowns in intimate conversations, taking into account the relationship context. Our approach draws on Nonviolent Communication (NVC) theory Rosenberg and Chopra (2015), a structured approach widely used in conflict resolution and therapeutic settings, to guide our evaluation of LLM’s detecting harmful communication. First, we introduce PersonaConflicts Corpus, a dataset of N=5,772N=5{,}772 simulated conflict dialogues between familiar social partners, spanning across diverse scenarios. For each conversation, we generate two distinct backstories: a positive backstory, which frames one character’s actions as more understandable or sympathetic, and a negative backstory, which portrays the same character in a more problematic or blameworthy light Moon et al. (2024). Through human validation, we show that scenarios and backstories are largely believable.
@@ 2.1 Conversational Breakdown Detection and Reframing
In the area of AIMC, text rewriting can be used to improve interpersonal outcomes like empathy or social connection by suggesting changes to the tone or style of a message at the right time Hancock et al. (2020). To support such systems, prior tasks propose detecting conversational breakdowns between people, and reframing messages to be more empathetic.
@@ 4 PersonaConflicts Corpus
We introduce the PersonaConflicts Corpus, a dataset of realistic conflict and non-conflict scenarios simulated using LLMs (Figure 3). Collecting large-scale real datasets of private, authentic breakdowns is non trivial, as these conversations are rarely shared publicly (unlike online or social media disputes), much less with backstories of the relationship history between individuals. Further, steps must be taken to ensure privacy, obtain consent, and address ethical considerations. As such, we draw on recent work in backstory generation and persona alignment Moon et al. (2024); Jiang et al. (2024); Hu and Collier (2024) as well as multi-agent simulation Zhou et al. (2023a, 2025); Ahmad et al. (2025); Kim et al. (2024) to generate conflict-laden scenarios which inject violent communication practices in turns and generate a set of non-conflict conversations. Our simulation setup uses the gpt-4 model and all prompts are included in Appendix A. We discuss our human evaluation verifying soundness of the dataset in Section 5.
@@ 5 Human Study and Annotation
To answer RQ1, how backstory influences people’s perceptions of conflict, we conducted a human study to assess the impact of backstory variant on perception of conflict at the conversation level and turn level, as well as to evaluate the quality of our dataset and obtain fine-grained annotations of violent or non-violent communication types. Our validation approach is grounded in prior work on synthetic dialogue evaluation Zhou et al. (2023a); Li et al. (2023); Zhan et al. (2023); Bao et al. (2023); Zhou et al. (2023b), which rely on human ratings of plausibility, naturalness, or coherence to validate generated conversations.
@@
We conducted a between-subjects study where participants are assigned to a positive or negative backstory. First, participants read the background of the characters and the conversation and rated overall measures of the conversation. Then, they were asked to rate each turn of the dialogue (See Figure 4 for an example of our user interface). The average work time was 17.28 minutes, and workers were paid \$3 for each HIT. Two independent workers completed each HIT for inter-annotator agreement calculation, resulting in a total of 480 annotations (3,474 turns rated across 120 conversations with 2 versions of backstory and 2 annotators per conversation). All annotation templates and discussion of quality controls are included in the Appendix.
@@
NEG Turn is problematic (4 point) .80.80 .44.44
Emotional impact (3 point) .79.79 .46.46
POS Turn is problematic (4 point) .78.78 .34.34
Emotional impact (3 point) .78.78 .42.42
@@ 6 Effect of Backstory Personalization on Human Participants
As shown in Figure 6, we found that providing relationship backstory significantly shifted participant perceptions of communication quality. Specifically, for negative backstories, characters were rated as more problematic both at the turn level ($t(1083)=3.73$, $p=0.0002$, Cohen’s $d=0.23$) and at the overall conversation level ($t(232)=4.18$, $p<0.0001$, Cohen’s $d=0.55$). Additionally, with negative backstories, participants found the character’s communication less understandable (t(232)=4.70t(232)=-4.70, $p<0.0001$, Cohen’s d=0.61d=-0.61) and interpreted their behavior as expressing more negative intent towards the other character (t(232)=4.95t(232)=-4.95, $p<0.0001$, Cohen’s d=0.65d=-0.65). Notably, participants expressed significantly lower sympathy towards the character when a negative backstory was present (t(232)=7.05t(232)=-7.05, $p<0.0001$, Cohen’s d=0.92d=-0.92), indicating a strong effect of relationship context on social judgments. These findings demonstrate that backstory personalization meaningfully influences how human raters interpret conflict.
@@
Table 2: F1 scores (in percentage) for predicting turn problematicness and emotional impact across models, conditions, and backstory types. Bold indicates the highest score in a model/backstory group; underline indicates the second highest. Significance stars denote statistical difference from the prior condition: * p<<0.05, ** p<<0.01, *** p<<0.001.
@@ 7.2 Results and Discussion
These results suggest that while GPT-4o is the most consistent with human perception across both tasks, LLaMA is prone to strong overestimation in both problematic detection and emotional response. Gemini’s behavior is more sensitive to backstory polarity, with notable shifts in prediction direction depending on whether a speaker is portrayed sympathetically or not, however these shifts are more extreme than human annotators’ ratings. Overall, our findings indicate that models might not be effectively leveraging relationship backstories to tailor understanding of conversational dynamics, in alignment with human perception. To further delve into these results, we include qualitative examples across models in Appendix E
@@ 8 Conclusion
Simulation-based data. Our dataset and evaluation pipeline offer a scalable and theory-informed way to study communication breakdowns, but the conversations are generated via simulation. Synthetic conversations may not fully capture the ecological validity of real-world interpersonal dynamics Wang et al. (2025), and the NVC-based four-step generation procedure may impose a more structured progression of conflict than naturally occurs. Language models can mirror patterns of speech and emotion, but they lack lived experience, embodied context, and nuanced power dynamics. Thus, findings from our simulated dialogue corpus may not fully generalize to naturally occurring conflicts, where nonverbal cues, cultural context, and relationship history shape interpretation in more complex ways. However, consistent with prior work in social dialogue simulation Bao et al. (2023); Chuang et al. (2024); Hu and Collier (2024); Zhou et al. (2023a); Li et al. (2023), our generated conversations were deemed largely naturalistic by human raters, and obtaining large-scale datasets of intimate, conflict-laden conversations between loved ones remains infeasible due to ethical and privacy constraints. We view our work as a proof-of-concept testbed rather than a replica of real-world phenomena, and future work should explore methods to bridge the gap between simulation and authentic data, such as incorporating real-world pilot studies or mixed human–synthetic evaluation designs Finch and Choi (2024).
Annotation and evaluation. Our validation relied on crowdsourced ratings of believability, following accepted practice in simulation-based dialogue studies. While we provided extensive guidelines and quality controls, believability captures whether a conversation could plausibly happen, not whether it is fully representative or authentic. Inter-rater agreement was moderate (Krippendorff’s $\alpha$ = 0.34–0.46), underscoring subjectivity in conflict judgments. We also focused our human study on a subset of key outcome measures (e.g., problematicness, sympathy, intention), leaving unexplored dimensions such as trust, perceived agency, or emotional volatility for future research. Automatic evaluation metrics on coherence and naturalness could also complement human ratings in future versions of the corpus.
Scope and generalizability. Our study is limited to single-modality (text) interactions, Western relationship contexts, and the English language. Cultural variation in conflict expression and resolution is well-documented Tschacher et al. (2014), and expanding to cross-cultural, multilingual, and multimodal settings (e.g., tone, pitch, and gestures) remains important. Moreover, our focus was on detection of conflict rather than modeling effective responses to conflict. While this narrower scope was intentional, we see our work as a first step toward contextualized conflict response generation in intimate interpersonal domains.
@@
All studies conducted in this work were classified under Institutional Review Board (IRB) exemption status. While our work aims to enhance interpersonal understanding and mitigate conflict through AIMC, the use of simulated dialogues and backstories about emotionally sensitive relationships—such as those between romantic partners or family members—raises concerns around realism and potential misuse. Although we do not collect or model real personal data, generated dialogues might still resemble real-life situations and emotional dynamics. If deployed in real-world applications, such systems could be used to influence perceptions of others, shape interpretations of interpersonal interactions, or even manipulate emotional outcomes, especially in high-stakes or abusive relationships. It is crucial that such tools remain assistive rather than prescriptive, providing support while preserving user autonomy and avoiding overreach in delicate relational contexts.
@@ Acknowledgements
We would like to thank all of our participants and teammates for their invaluable contributions to this project. Special thanks to Ashish Sharma and Shannon Shen for feedback on the project. This work was funded in part by DSO National Laboratories and supported by the Defense Advanced Research Projects Agency (DARPA) under Agreement No. HR00112490410.
@@ A.5 Model Experiment Prompts
Appendix B MTurk Quality Controls
To ensure reliable annotations, we recruited experienced MTurk workers with Master’s status, implemented attention checks, filtered low-effort responses, and enforced minimum completion times. Prior studies of interpersonal conflict and toxicity (e.g., ToxiGen Hartvigsen et al. (2022), Unintended Offense Tsai et al. (2024), Sotopia Zhou et al. (2023a)) similarly rely on diverse crowdworkers, who bring lived social experience to interpreting interpersonal exchanges.
Appendix C Annotation Templates
Appendix D Supplementary Results
@@
Appendix E Qualitative Examples of Model Error
GPT-4o example for positive backstory: Jaxon was the eldest amongst his siblings and has always been the protector, especially of Emily, who was the youngest. Even during their childhood days, Jaxon played the role of a gentle critic, pointing out Emily’s potential growth areas in her drawings and writings, nurturing her ability to improve and perform better. Their parents often commended him for his supportive and yet constructive attitude towards Emily. And it was this constant guidance and criticism from Jaxon that played a significant role in Emily developing her writing skills, leading her to become a recognized writer under a pseudonym. When Jaxon came across Emily’s secret, he couldn’t help but continue his role as a supportive critic, hoping to elevate her writing to even greater heights, just as he had during their childhood.
      Turn: "(Jaxon): Hey Em, I read one of your latest novels and..."    Human Rating: 1.5 (slightly negative feeling)    Model Rating: 3.0 (positive feeling)
Llama example for negative backstory: Baxter has always been a loner and finds comfort in the isolation of his work. His lack of social interaction led him to develop a cynical attitude towards people and their capacities. He’d been friends with Isabelle’s parents and saw Isabelle grow up. When Isabelle’s parents passed away in an accident, Baxter - unable to process his grief - began to push her away, masking his fears behind a facade of humor and sarcasm. Still considering Isabelle as the young, naive girl he once knew, his jokes became more demeaning over time, depreciating her efforts in her job and personal growth. His remarks started to create a distance between them, leading to an increasingly toxic relationship.
      Turn: "Isabelle: You know what, Baxter? I’ve been trying to ignore your snarky comments for a while, but this is starting to hurt."    Human Rating: 1.0 (not really problematic)    Model Rating: 3.0 (moderately problematic)
Llama example for positive backstory: Leo had years of experience juggling between his work as a dentist and taking care of his beloved daughter. When his daughter’s mother left, he found himself thrust into a dual role of being both a father and a mother. Dealing with this dynamic was challenging yet it taught Leo the value of family and instilled in him a deep sense of responsibility. As the elder brother to Naomi, he felt it was his duty to ensure he was there for her, just as he was for his little daughter. Having lost their own parents at an early age, Leo was adamant in providing a strong family base for his sister. Naomi was more than just Leo’s sister, she was also his confidante and friend. Leo’s unwavering dedication to family often led him to bear the responsibility of solving everything on his own.
      Turn: "(Leo): I think I’ve been so tied up recently with work...sometimes I overlook things. Family stuff, you know?"    Human Rating: 2.0 (neutral feeling)    Model Rating: 3.0 (positive feeling)
Gemini examples for negative backstory: Baxter has always been a loner and finds comfort in the isolation of his work. His lack of social interaction led him to develop a cynical attitude towards people and their capacities. He’d been friends with Isabelle’s parents and saw Isabelle grow up. When Isabelle’s parents passed away in an accident, Baxter - unable to process his grief - began to push her away, masking his fears behind a facade of humor and sarcasm. Still considering Isabelle as the young, naive girl he once knew, his jokes became more demeaning over time, depreciating her efforts in her job and personal growth. His remarks started to create a distance between them, leading to an increasingly toxic relationship.
      Turn: "Baxter: You’re right... I’m sorry. I shouldn’t have said those things."    Human Rating: 1.0 (not really problematic)    Model Rating: 3.0 (moderately problematic)
      Turn: "Baxter: Small? I’ve seen you handle customers double your age with that articulation of yours."    Human Rating: 2.5 (moderately positive feeling)    Model Rating: 1.0 (negative feeling)